# Geometric and topological rigidity of pinched submanifolds II

### v1

@@ 3 Proofs of the main results
Let $f\colon M^{n}\to\mathbb{Q}_{c}^{n+m},n\geq 4,c\geq 0$, be a substantial isometric immersion of a compact, simply connected, even-dimensional manifold with flat normal bundle. Assume that equality holds in (italic-$\ast$1) for $k=n/2$ at every point, and that there exist two principal normals of multiplicity $k$. Then, $M^{n}$ is isometric to a torus $\mathbb{S}^{k}(r)\times\mathbb{S}^{k}(\sqrt{R^{2}-r^{2}})$ and $f$ is a composition $f=j\circ g$, where $g\colon M^{n}\to\mathbb{S}^{n+1}(R)$ is the standard embedding of the torus $\mathbb{S}^{k}(r)\times\mathbb{S}^{k}(\sqrt{R^{2}-r^{2}})$ into a sphere $\mathbb{S}^{n+1}(R)$, and $j\colon\mathbb{S}^{n+1}(R)\to\mathbb{Q}_{c}^{n+2}$ is an umbilical inclusion, with $R\leq 1$ if $c=1$.
@@
Let $\eta_{1},\eta_{2}$ be the principal normals with corresponding distributions $E_{1},E_{2}$. Since $2\mathcal{H}_{f}=\eta_{1}+\eta_{2}$ and $S=k(\|\eta_{1}\|^{2}+\|\eta_{2}\|^{2})$, condition ($\ast$1), which now holds as equality for $k=n/2$, implies
@@
Consider the open subset $M^{*}$ of points where $f$ is not umbilical, i.e., where $\eta_{1}\neq\eta_{2}$. On $M^{*}$, the vector fields $\eta_{1}$ and $\eta_{2}$ are smooth Dupin principal normals, with smooth corresponding distributions $E_{1},E_{2}$, each of rank $k$.
@@
for all $X,Y\in E_{i},Z\in E_{j},1\leq i,j\leq 2$. Using (13) and the fact that $\eta_{1}$ and $\eta_{2}$ are Dupin principal normal vector fields, (14) implies that the following relations hold on $M^{*}$:
@@
Case $c>0$. From (13), it is clear that $M^{*}=M^{n}$. Consequently, it follows from (15) that the distributions $E_{1}$ and $E_{2}$ are totally geodesic on $M^{n}$. Since $M^{n}$ is simply connected, it is well-known that $M^{n}$ is a Riemannian product $M_{1}^{k}\times M_{2}^{k}$ (cf. Theorem $8.2$ in [5]) such that $TM_{i}^{k}=E_{i}$, $i=1,2$. As the second fundamental form of $f$ is adapted to the distributions $E_{1}$ and $E_{2}$, then Theorem $8.4$ and Corollary $8.6$ in [5] imply that the submanifold is an extrinsic product of isometric immersions, each of which is totally umbilical. Therefore, the submanifold is a torus $\mathbb{S}^{k}(r)\times\mathbb{S}^{k}(\sqrt{R^{2}-r^{2}})$ in a sphere $\mathbb{S}^{n+1}(R)\subset\mathbb{S}^{n+2}$.
@@
Now, suppose that the interior ${\rm int}(M_{0})$ of the subset $M_{0}=M^{*}\smallsetminus M_{+}$ is nonempty. Assume that $\eta_{2}=0$ on a connected component $U$ of ${\rm int}(M_{0})$. Then $f$ has constant index of relative nullity $k$ on $U$. From (15), it follows that the relative nullity distribution $\mathcal{D}_{f}=E_{2}$ is totally geodesic. Let $C_{T}\colon\mathcal{D}_{f}^{\perp}\to\mathcal{D}_{f}^{\perp}$ be the associated splitting tensor for any $T\in\mathcal{D}_{f}$ (see [5, p. 186]). Since
@@
$(\nabla_{X}^{\perp}\alpha_{f})(Y,T)=(\nabla_{T}^{\perp}\alpha_{f})(X,Y)\;\;{\text{for all}}\;\;X,Y\in\mathcal{D}_{f}^{\perp},T\in\mathcal{D}_{f}$
implies that $C_{T}={\langle}\nabla\log\|\eta_{1}\|,T{\rangle}Id$, where $Id$ is the identity map on the conullity distribution $\mathcal{D}_{f}^{\perp}$. This shows that the conullity distribution is umbilical, and consequently, integrable.
@@
Thus, $f\circ j_{\varSigma}$ is umbilical and $g(\varSigma^{k})$ is a sphere $\mathbb{S}^{k}(r)$ centered at a point $x_{0}\in\mathbb{R}^{n+m}$ with radius $r=1/\sqrt{\lambda^{2}+\|\nabla\log\|\eta_{1}\|\|^{2}}$. If the umbilical submanifold $\mathbb{Q}_{\tilde{c}}^{k+m}$ is totally geodesic in $\mathbb{R}^{n+m}$, then the submanifold $f$ is a $k$-cylinder over the sphere $\mathbb{S}^{k}(r)$. Alternatively, if $\mathbb{Q}_{\tilde{c}}^{k+m}$ is a sphere centered at a point $\tilde{x}_{0}\in\mathbb{R}^{n+m}$, then $f$ is a $(k-1)$-cylinder over a submanifold $N^{k+1}$, which is a cone over the sphere $\mathbb{S}^{k}(r)$ with its vertex at $\tilde{x}_{0}\neq x_{0}$.
Consequently, on ${\rm int}(M_{0})$ the submanifold is locally a $k$-cylinder in $\mathbb{R}^{n+1}$ over a sphere $\mathbb{S}^{k}(r)$, or a $(k-1)$-cylinder over a submanifold $N^{k+1}$ in $\mathbb{R}^{k+2}$, which is a cone over a sphere $\mathbb{S}^{k}(r)$.
In the later case, $f$ is given locally on $\mathrm{int}(M_{0})$ by
@@
where $\mathbb{S}^{k+m}(R)\subset\mathbb{R}^{k+m+1}$, $\mathbb{R}^{n+m}=\mathbb{R}^{k+m+1}\oplus\mathbb{R}^{k-1}$, and $\{v_{\ell}\}_{1\leq\ell\leq k-1}$ is an orthonormal basis of $\mathbb{R}^{k-1}$. The Laplacian operator $\Delta_{M}$ of $M^{n}$ is given by
@@
Consequently, either $\Delta_{M}S\geq 0$ or $\Delta_{M}S\leq 0$ on ${\rm int}(M_{0})$, depending on the dimension $n$.
@@
Thus, on $M_{0}$, one of the Dupin principal normal vector fields vanishes, while the other has constant length $S/k$. On the other hand, both principal normals have constant and positive length on each connected component of $M_{+}$. By continuity and the compactness of $M^{n}$, it follows that $M_{+}=M^{n}$. Consequently, using a similar argument as in Case $c>0$, we conclude that $M^{n}$ is isometric to a torus $\mathbb{S}^{k}(r)\times\mathbb{S}^{k}(\sqrt{R^{2}-r^{2}})$, and $f$ is the standard embedding into a sphere $\mathbb{S}^{n+1}(R)\subset\mathbb{R}^{n+2}$. ∎
Proof of Theorem 2: Part $(i)$ of Proposition 8 implies that $M^{4}$ has nonnegative isotropic curvature and that the Bochner-Weitzenböck operator is nonnegative at any point. The proof of the theorem is divided into three cases.
@@
Since $M^{4}$ is homeomorphic to $\mathbb{S}^{4}$, case (ii) above is ruled out. Clearly, in case (i), the manifold must be diffeomorphic to $\mathbb{S}^{4}$. Furthermore, the only 4-dimensional compact symmetric spaces are the round spheres, the product of two 2-dimensional spheres, or the complex or the quaternionic projective space $\mathbb{H}P^{1}$. Consequently, in case (iii), the manifold $M^{4}$ must be diffeomorphic to $\mathbb{S}^{4}$.
@@
Case (b). Since $M^{4}$ is simply connected, the surface $\varSigma^{2}$ must be diffeomorphic to $\mathbb{S}^{2}$, implying that $M^{4}$ is diffeomorphic to $\mathbb{S}^{2}\times\mathbb{S}^{2}$. Consequently, $\beta_{\pm}(M)=1$. It then follows from part $(ii)$ of Proposition 8 that equality holds in ($\ast$1). Moreover, at each point $x\in M^{4}$, there exists an oriented orthonormal four-frame $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{x}M$ such that the second fundamental form $\alpha_{f}$ of $f$ satisfies the conditions (3)-(5).
In addition, there exist a nontrivial self-dual harmonic 2-form $\omega_{+}$ and a nontrivial anti-self-dual harmonic 2-form $\omega_{-}$. By the Bochner-Weitzenböck formula, both $\omega_{+}$ and $\omega_{-}$ are parallel and ${\langle}\mathcal{B}^{[2]}(\omega_{\pm}),\omega_{\pm}{\rangle}=0$ at any point. Consequently, $\ker\mathcal{B}_{+}^{[2]}\neq 0$ and $\ker\mathcal{B}_{-}^{[2]}\neq 0$ at any point. Using Lemma (9) and (3)-(5), we deduce that only cases $(ii1)$ and $(iii1)$ in that lemma can occur at any point. Furthermore, applying (3)-(5) once more, it follows that four-frame $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{x}M$ diagonalizes the second fundamental form at any $x\in M^{4}$. As a result, $f$ has flat normal bundle. Moreover, $\alpha_{11}$ and $\alpha_{44}$ are principal normals of $f$ with multiplicity $2$ satisfying ${\langle}\alpha_{11},\alpha_{44}{\rangle}=-c$. It then it follows from Proposition 10 that the submanifold is as described in part $(iiia)$ of the theorem.
@@
Case (c). In this case, either $\beta_{+}(M)=1$ and $\beta_{-}(M)=0$, or $\beta_{+}(M)=0$ and $\beta_{-}(M)=1$. We will only treat the former case, as the latter one can be handled in a similar manner. It follows then from part $(ii)$ of Proposition 8 that equality holds in ($\ast$1). Furthermore, at each point $x\in M^{4}$ there exists an oriented orthonormal four-frame $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{x}M$ such that the second fundamental form $\alpha_{f}$ of $f$ satisfies the conditions (3)-(5).
@@
Clearly, there exists a nontrivial self-dual harmonic 2-form $\omega_{+}$. By the Bochner-Weitzenböck formula, the form $\omega_{+}$ is parallel, and ${\langle}\mathcal{B}^{[2]}(\omega_{+}),\omega_{+}{\rangle}=0$ at any point. Hence, $\ker\mathcal{B}_{+}^{[2]}\neq 0$ at any point, and part $(ii)$ of Lemma (9) applies. Using (3) and (5), we conclude that only case $(ii1)$ in that lemma can occur at any point. Thus, we have
$\alpha_{14}+\alpha_{23}=\alpha_{13}-\alpha_{24}=0.$ (16)
We assert that $\mu^{+}_{2}\mu^{+}_{3}>0$ at some point. Since $c\geq 0$, if $\mu^{+}_{2}\mu^{+}_{3}=0$ at a point $x$, it follows that $c=0$. Using (3)-(5), we deduce that $f$ must be totally geodesic at $x$. Therefore, there must exist a point $x_{0}$ where $\mu^{+}_{2}\mu^{+}_{3}>0$. From part $(i)$ of Lemma 9, it follows that the kernel of $\mathcal{B}_{+}^{[2]}$ at $x_{0}$ is spanned by the vector $e_{1}\wedge e_{2}+e_{3}\wedge e_{4}$.
@@
Let $X$ be the dual to the self-dual form $\omega_{+}$. Since this form is parallel, we may assume, after possibly multiplying by a constant such that $\|\omega_{+}\|=\sqrt{2}$, that at the point $x_{0}$ we have $X_{x_{0}}=e_{1}\wedge e_{2}+e_{3}\wedge e_{4}.$ Now, consider the almost complex structure $J_{x_{0}}\colon T_{x_{0}}M\to T_{x_{0}}M$ given by $J_{x_{0}}e_{1}=e_{2}$ and $J_{x_{0}}e_{3}=e_{4}$. Then, we have
$\omega_{+}(v,w)={\langle}v,J_{x_{0}}w{\rangle}\;\;\text{for any}\>\>v,w\in T_{x_{0}}M.$
@@
$\omega_{+}(X,Y)={\langle}X,JY{\rangle}\;\;\text{for any}\>\>X,Y\in\mathcal{X}(M).$
Clearly, $J$ is parallel because $\omega_{+}$ is parallel. Now we claim that $J$ is orthogonal, i.e., $\|J_{x}v\|=\|v\|$ for any point $x\in M^{4}$ and any $v\in T_{x}M$. Indeed, let $V$ be a parallel vector field along a curve $c\colon[0,1]\to M$ such that $c(0)=x,c(1)=x_{0}$ and $V(0)=v$. Obviously, the vector field $W=JV$ is also parallel along $c$. Using that $J_{x_{0}}$ is orthogonal we have
@@
which proves the claim. Since $J$ is both skew-symmetric and orthogonal, we have that $J$ is an almost complex structure that is also parallel. Hence, the triple $(M^{4},{\langle}\cdot,\cdot{\rangle},J)$ is a Kähler manifold.
In case $c>0$, we have that $\mu_{1}^{+}=0$ and $\mu_{2}^{+}\mu_{3}^{+}>0$ at any point. Hence, $X_{x}=e_{1}\wedge e_{2}+e_{3}\wedge e_{4}$ at any point $x$. Thus, we have that $Je_{1}=e_{2}$ and $Je_{3}=e_{4}$ at any point. From (16), it follows that the second fundamental form of the submanifold satisfies
@@
In case $c=0$, the above argument applies to the open subset of points where $f$ is not totally geodesic. By continuity, (17) holds everywhere.
Immersions satisfying condition (17) have parallel second fundamental form (see [7, Th. 4]). Moreover, under the immersion $f$, each geodesic of $M$ is mapped into a plane circle. Such submanifolds were classified in [7] and [18]. From this classification, it follows that the submanifold is as described in part $(iiib)$ of the theorem, thereby completing the proof for the case where the manifold is simply connected.

### v2

@@ 3 Proofs of the main results
Let $f\colon M^{n}\to\mathbb{Q}_{c}^{n+m},n\geq 4,c\geq 0$, be a substantial isometric immersion of a compact, simply connected, even-dimensional manifold with flat normal bundle. Assume that equality holds in ($\ast$1) for $k=n/2$, and that there exist two distinct Dupin principal normals $\eta_{1}$ and $\eta_{2}$, each of multiplicity $k$, at every point $x\in M^{n}$ where $f$ is not umbilical, such that
$T_{x}M=E_{\eta_{1}}(x)\oplus E_{\eta_{2}}(x).$
Then, $M^{n}$ is isometric to a torus $\mathbb{S}^{k}(r)\times\mathbb{S}^{k}(\sqrt{R^{2}-r^{2}})$ and $f$ is a composition $f=j\circ g$, where $g\colon M^{n}\to\mathbb{S}^{n+1}(R)$ is the standard embedding of the torus $\mathbb{S}^{k}(r)\times\mathbb{S}^{k}(\sqrt{R^{2}-r^{2}})$ into a sphere $\mathbb{S}^{n+1}(R)$, and $j\colon\mathbb{S}^{n+1}(R)\to\mathbb{Q}_{c}^{n+2}$ is an umbilical inclusion, with $R\leq 1$ if $c=1$.
@@
Consider the open subset $M^{*}$ of points where $f$ is not umbilical. On $M^{*}$, the vector fields $\eta_{1}$ and $\eta_{2}$ are smooth Dupin principal normals, with smooth corresponding distributions $E_{1}=E_{\eta_{1}}$ and $E_{2}=E_{\eta_{2}}$, each of rank $k$ (see Proposition 1 in [21] or Lemma 10 in [28]).
Since
$2\mathcal{H}_{f}=\eta_{1}+\eta_{2}\;\;{\text{and}}\;\,S=k(\|\eta_{1}\|^{2}+\|\eta_{2}\|^{2}),$
condition ($\ast$1), which now holds as equality for $k=n/2$, implies
@@
for all $X,Y\in E_{i},Z\in E_{j},1\leq i,j\leq 2$. Using equation (14) and the fact that the Dupin principal normal vector fields $\eta_{1}$ and $\eta_{2}$ are parallel in the normal connection along $E_{1}$ and $E_{2}$, respectively (see Proposition 1.22 in [5]), it follows from equation (15) that the following relations hold on $M^{*}$:
@@
Case $c>0$. From equation (14), it follows that $M^{*}=M^{n}$. Consequently, equation (16) implies that the distributions $E_{1}$ and $E_{2}$ are totally geodesic on $M^{n}$. Since $M^{n}$ is simply connected, it is well-known that $M^{n}$ is a Riemannian product $M_{1}^{k}\times M_{2}^{k}$ (cf. Theorem $8.2$ in [5]) such that $TM_{i}^{k}=E_{i}$, $i=1,2$. As the second fundamental form of $f$ is adapted to the distributions $E_{1}$ and $E_{2}$, then Theorem $8.4$ and Corollary $8.6$ in [5] imply that the submanifold is an extrinsic product of isometric immersions, each of which is totally umbilical. Therefore, the submanifold is a torus $\mathbb{S}^{k}(r)\times\mathbb{S}^{k}(\sqrt{R^{2}-r^{2}})$ in a sphere $\mathbb{S}^{n+1}(R)\subset\mathbb{S}^{n+2}$.
@@
Now, suppose that the interior $\operatorname{int}(M_{0})$ of the subset $M_{0}=M^{*}\smallsetminus M_{+}$ is nonempty. Assume that $\eta_{2}=0$ on a connected component $U$ of $\operatorname{int}(M_{0})$. Then $f$ has constant index of relative nullity $k$ on $U$. From (16), it follows that the relative nullity distribution $\mathcal{D}_{f}=E_{2}$ is totally geodesic. Let $C_{T}\colon\mathcal{D}_{f}^{\perp}\to\mathcal{D}_{f}^{\perp}$ be the associated splitting tensor for any $T\in\mathcal{D}_{f}$ (see [5, p. 186]). Since
@@
[…]
implies that
$C_{T}={\langle}\nabla\log\|\eta_{1}\|,T{\rangle}Id,$
where $Id$ is the identity map on the conullity distribution $\mathcal{D}_{f}^{\perp}$. This shows that the conullity distribution is umbilical, and consequently, integrable.
@@
Thus, $f\circ j_{\varSigma}$ is umbilical and $g(\varSigma^{k})$ is a sphere $\mathbb{S}^{k}(r)$ centered at a point $x_{0}\in\mathbb{R}^{n+m}$ with radius
$r=\frac{1}{\sqrt{\|\eta_{1}\|^{2}+\|\nabla\log\|\eta_{1}\|\|^{2}}}.$
If the umbilical submanifold $\mathbb{Q}_{\tilde{c}}^{k+m}$ is totally geodesic in $\mathbb{R}^{n+m}$, then the submanifold $f$ is a $k$-cylinder over the sphere $\mathbb{S}^{k}(r)$. Alternatively, if $\mathbb{Q}_{\tilde{c}}^{k+m}$ is a sphere centered at a point $\tilde{x}_{0}\in\mathbb{R}^{n+m}$, then $f$ is a (k1)(k-1)-cylinder over a submanifold $N^{k+1}$, which is a cone over the sphere $\mathbb{S}^{k}(r)$ with its vertex at $\tilde{x}_{0}\neq x_{0}$.
Consequently, on $\operatorname{int}(M_{0})$ the submanifold is locally a $k$-cylinder in $\mathbb{R}^{n+1}$ over a sphere $\mathbb{S}^{k}(r)$, or a (k1)(k-1)-cylinder over a submanifold $N^{k+1}$ in $\mathbb{R}^{k+2}$, which is a cone over a sphere $\mathbb{S}^{k}(r)$.
In the later case, $f$ is given locally on $\operatorname{int}(M_{0})$ by
@@
where
$\mathbb{S}^{k+m}(R)\subset\mathbb{R}^{k+m+1},\;\;\mathbb{R}^{n+m}=\mathbb{R}^{k+m+1}\oplus\mathbb{R}^{k-1},$
and $\{v_{\ell}\}_{1\leq\ell\leq k-1}$ is an orthonormal basis of $\mathbb{R}^{k-1}$. The Laplacian operator $\Delta_{M}$ of $M^{n}$ is given by
@@
Consequently, either $\Delta_{M}S\geq 0$ or $\Delta_{M}S\leq 0$ on $\operatorname{int}(M_{0})$, depending on the dimension $n$.
@@
Thus, on $M_{0}$, one of the Dupin principal normal vector fields vanishes, while the other has constant length $S/k$. On the other hand, both principal normal fields have constant and positive length on each connected component of $M_{+}$. By continuity and the compactness of $M^{n}$, it follows that $M_{+}=M^{n}$. Consequently, using a similar argument as in Case $c>0$, we conclude that $M^{n}$ is isometric to a torus $\mathbb{S}^{k}(r)\times\mathbb{S}^{k}(\sqrt{R^{2}-r^{2}})$, and the immersion $f$ is the standard embedding into a sphere $\mathbb{S}^{n+1}(R)\subset\mathbb{R}^{n+2}$. ∎
Proof of Theorem 2: Part $(i)$ of Proposition 8 implies that $M^{4}$ has nonnegative isotropic curvature and that the Bochner-Weitzenböck operator $\mathcal{B}^{[2]}$ is nonnegative at every point. The proof of the theorem is divided into three cases.
@@
Since $M^{4}$ is homeomorphic to $\mathbb{S}^{4}$, case (ii) above is excluded. In case (i), the manifold is clearly diffeomorphic to $\mathbb{S}^{4}$. Moreover, the only compact symmetric spaces of dimension four are the round sphere, the product of two 2-dimensional spheres, the complex projective plane, and the quaternionic projective line $\mathbb{H}P^{1}\cong\mathbb{S}^{4}$. Consequently, in case (iii), the manifold $M^{4}$ must also be diffeomorphic to $\mathbb{S}^{4}$.
@@
Case (b). Since $M^{4}$ is simply connected, the surface $\varSigma^{2}$ must be diffeomorphic to $\mathbb{S}^{2}$, which implies that $M^{4}$ is diffeomorphic to $\mathbb{S}^{2}\times\mathbb{S}^{2}$. Consequently, $\beta_{\pm}(M)=1$. It then follows from part $(ii)$ of Proposition 8 that equality holds in ($\ast$1). Moreover, at each point $x\in M^{4}$, there exists an oriented orthonormal four-frame $\{e_{1},e_{2},e_{3},e_{4}\}\subset T_{x}M$ such that the second fundamental form $\alpha_{f}$ of $f$ satisfies conditions (3)-(5).
In addition, there exist a nontrivial self-dual harmonic 2-form $\omega_{+}$ and a nontrivial anti-self-dual harmonic 2-form $\omega_{-}$. By the Bochner-Weitzenböck formula, both $\omega_{+}$ and $\omega_{-}$ are parallel and
${\langle}\mathcal{B}^{[2]}(\omega_{\pm}),\omega_{\pm}{\rangle}=0$
everywhere. Consequently, $\ker\mathcal{B}_{+}^{[2]}\neq 0$ and $\ker\mathcal{B}_{-}^{[2]}\neq 0$ at every point. Hence, both parts $(ii)$ and $(iii)$ of Lemma 9 hold everywhere.
We claim that the immersion $f$ has flat normal bundle at every point $x\in M^{4}$. Suppose first that both cases (ii1) and (iii1) hold simultaneously at $x$. Then, it follows from (3) and (5) that $c=0$, and hence $f$ is totally geodesic at this point. If, instead, either (ii1) or (iii1) does not hold at $x$, then Lemma 9 implies that $f$ has flat normal bundle at this point. This proves the claim that $f$ has flat normal bundle.
Since by part $(ii)$ of Proposition 8, the shape operator, associated with any normal direction, has at most two distinct eigenvalues, each of multiplicity $2$ at every point. This together with the flatness of the normal bundle implies that there exist two distinct Dupin principal normals, $\eta_{1}$ and $\eta_{2}$ each of multiplicity $2$, at every point $x\in M^{4}$ where $f$ is not umbilical, such that
$T_{x}M=E_{\eta_{1}}(x)\oplus E_{\eta_{2}}(x).$
It then follows from Proposition 10 that the submanifold is as described in part $(iii$-$a)$ of the theorem.
Case (c). In this case, either
$\beta_{+}(M)=1\;\;\text{and}\;\;\beta_{-}(M)=0,\;\;\text{or}\;\;\beta_{+}(M)=0\;\;\text{and}\;\;\beta_{-}(M)=1.$
@@
We will treat only the former case, as the latter can be handled in a similar manner. It then follows from part $(ii)$ of Proposition 8 that equality holds in ($\ast$1). Moreover, at each point $x\in M^{4}$, there exists an oriented orthonormal four-frame $\{e^{x}_{1},e^{x}_{2},e^{x}_{3},e^{x}_{4}\}\subset T_{x}M$ such that the second fundamental form $\alpha_{f}$ of $f$ satisfies conditions (3)-(5).
@@
Clearly, there exists a nontrivial self-dual harmonic 2-form $\omega_{+}$. By the Bochner-Weitzenböck formula, the form $\omega_{+}$ is parallel, and
${\langle}\mathcal{B}^{[2]}(\omega_{+}),\omega_{+}{\rangle}=0$
at every point. Hence, $\ker\mathcal{B}_{+}^{[2]}\neq 0$ at every point, and part $(ii)$ of Lemma 9 applies.
We consider the subset
$M_{1}=\left\{x\in M^{4}:\mu^{+}_{1}(x)=0\;\;\text{and}\;\;\mu^{+}_{2}(x)\mu^{+}_{3}(x)>0\right\},$
where $\mu^{+}_{1},\mu^{+}_{2}$, and $\mu^{+}_{3}$ are as given in Lemma 9. It follows that
$M_{1}=\left\{x\in M^{4}:\mu^{+}_{1}(x)=0\;\;\text{and}\;\;\|\alpha_{11}-\alpha_{44}\|^{2}+4\|\alpha_{13}\|^{2}+4\|\alpha_{14}\|^{2}>0\right\}.$
We claim that this subset is nonempty. Suppose, to the contrary, that $M_{1}$ is empty. Then, at every point of $M^{4}$, one of the following conditions must hold:
$\alpha_{14}+\alpha_{23}\neq 0,\;\text{or}\;\alpha_{13}-\alpha_{24}\neq 0,\;\text{or}\;\|\alpha_{11}-\alpha_{44}\|^{2}+4\|\alpha_{13}\|^{2}+4\|\alpha_{14}\|^{2}=0.$ (17)
We claim that the immersion $f$ has flat normal bundle. Indeed, if the third condition in (17) holds at a point $x$, then (3) and (4) imply that $c=0$, and hence $f$ is totally geodesic at $x$. If, instead, either the first or the second condition in (17) holds at a point $x$, then part $(ii)$ of Lemma 9 shows that either case (ii2) or case (iii2) holds at that point. In either situation, Lemma 9 further implies that the submanifold $f$ has flat normal bundle and satisfies
$\dim N_{1}f\leq 2$
at all points of $M^{4}$.
By part $(ii)$ of Proposition 8, the shape operator, associated with any normal direction, has at most two distinct eigenvalues, each of multiplicity $2$ at every point. Combined with the flatness of the normal bundle, this implies the existence of two distinct Dupin principal normals, $\eta_{1}$ and $\eta_{2}$, each of multiplicity $2$, at every point $x\in M$ where $f$ is not umbilical, such that
$T_{x}M=E_{\eta_{1}}(x)\oplus E_{\eta_{2}}(x).$
Since the submanifold is substantial, Proposition 10 implies that its codimension is $m=2$. Consequently, the complex projective plane $\mathord{\mathbb{C}}\mathord{\mathbb{P}}^{2}$ would admit an immersion into $\mathbb{R}^{6}$ or in $\mathbb{S}^{6}$. This, however, contradicts known topological obstructions (see, for instance, [7, pp. 290-291]). Hence, $M_{1}$ must be nonempty.
From part $(i)$ of Lemma 9, it follows that at every point of $M_{1}$ we have
$\mu_{1}^{+}=0,\;\;\mu_{2}^{+}=\mu_{3}^{+}>0.$
Consequently, the kernel of $\mathcal{B}_{+}^{[2]}$ at each point $x\in M_{1}$ is spanned by the vector $e^{x}_{1}\wedge e^{x}_{2}+e^{x}_{3}\wedge e^{x}_{4}$.
@@
Let $Z$ be the dual to the self-dual form $\omega_{+}$. Since this form is parallel, we may normalize it (after multiplying by a constant if necessary) so that $\|\omega_{+}\|=\sqrt{2}$. Fix a point $x_{0}\in M_{1}$. At this point, we may assume that
$Z_{x_{0}}=e^{x_{0}}_{1}\wedge e^{x_{0}}_{2}+e^{x_{0}}_{3}\wedge e^{x_{0}}_{4}.$
Now, consider the almost complex structure $J_{x_{0}}\colon T_{x_{0}}M\to T_{x_{0}}M$ given by $J_{x_{0}}e^{x_{0}}_{1}=e^{x_{0}}_{2}$ and $J_{x_{0}}e^{x_{0}}_{3}=e^{x_{0}}_{4}$. Then, we have
$\omega_{+}(v,w)={\langle}J_{x_{0}}v,w{\rangle}\;\;\text{for all}\;\,v,w\in T_{x_{0}}M.$
@@
$\omega_{+}(X,Y)={\langle}JX,Y{\rangle}\;\;\text{for all}\;\,X,Y\in\mathcal{X}(M).$
Clearly, $J$ is parallel because $\omega_{+}$ is parallel. We claim that $J$ is orthogonal; that is, $\|J_{x}v\|=\|v\|$ for every point $x\in M^{4}$ and every $v\in T_{x}M$. Indeed, let $V$ be a parallel vector field along a curve $c\colon[0,1]\to M$ such that $c(0)=x,c(1)=x_{0}$ and $V(0)=v$. Obviously, the vector field $W=JV$ is also parallel along $c$. Using that $J_{x_{0}}$ is orthogonal we have
@@
which proves the claim. Since $J$ is both skew-symmetric and orthogonal, it defines an almost complex structure that is also parallel. Hence, the triple $(M^{4},{\langle}\cdot,\cdot{\rangle},J)$ is a Kähler manifold.
We now claim that, at every point $x\in M^{4}$, we have
$\alpha_{f}(J_{x}v,J_{x}w)=\alpha_{f}(v,w)\;\;\text{for all}\;\;v,w\in T_{x}M.$ (18)
To prove this, we distinguish two cases.
First, let $x$ be an arbitrary point in $M_{1}$. Since the kernel of $\mathcal{B}_{+}^{[2]}$ at $x$ is spanned by the vector $e^{x}_{1}\wedge e^{x}_{2}+e^{x}_{3}\wedge e^{x}_{4}$, and since $\|Z_{x}\|=\sqrt{2}$, it follows that
$Z_{x}=\pm\left(e^{x}_{1}\wedge e^{x}_{2}+e^{x}_{3}\wedge e^{x}_{4}\right).$
Because $Z$ is the dual of the self-dual form $\omega_{+}$, we have
${\langle}J_{x}v,w{\rangle}={\langle}{\langle}Z_{x},v\wedge w{\rangle}{\rangle}\;\;\text{for all}\;\,v,w\in T_{x}M.$
Consequently,
$J_{x}e^{x}_{1}=\pm e^{x}_{2}\;\;\text{and}\;\;J_{x}e^{x}_{3}=\pm e^{x}_{4}$
at each point $x\in M_{1}$. On the other hand, we have
$\alpha_{14}+\alpha_{23}=\alpha_{13}-\alpha_{24}=0$ (19)
at every point of $M_{1}$. Combining (3) with (19), we obtain
$\alpha_{f}(J_{x}e^{x}_{i},J_{x}e^{x}_{j})=\alpha_{ij}\;\;\text{for all}\;\;1\leq i,j\leq 4,$ (20)
and thus (18) follows from (20) by linearity.
Having analyzed the case $x\in M_{1}$, we now turn to the complementary case $x\notin M_{1}$. In this case, (17) holds at $x$.
If the third condition in (17) is satisfied, then (3) and (5) imply that $c=0$, and hence $f$ is totally geodesic at $x$. In this situation, (18) holds trivially.
Next, suppose that the first or the second condition in (17) holds at $x$. We treat only the first case, since the second can be handled in a similar way, and will therefore be omitted. Then part $(ii)$ of Lemma 9 shows that case (ii2) holds at that point. Together with (ii2), it follows from part $(i)$ of this lemma that the matrix of the Bochner-Weitzenböck operator $\mathcal{B}_{+}^{[2]}$ at $x$, with respect to the basis $\{\eta_{i}\}_{1\leq i\leq 3}$ of $\Lambda_{+}^{2}T_{x}M$, is given by
$4\|\alpha_{14}\|^{2}\begin{bmatrix}1&\rho&0\\ \rho&\rho^{2}&0\\ 0&0&\rho^{2}+1&\!\!\!\!\!\end{bmatrix}$,
where $\rho$ is the constant appearing in case (ii2) of Lemma 9. Therefore, the kernel of $\mathcal{B}_{+}^{[2]}$ at $x$ is spanned by the vector $\rho\eta_{1}+\eta_{2}$, and since $\|Z_{x}\|=\sqrt{2}$, it follows that
$Z_{x}=\lambda\left(-\rho\eta_{1}+\eta_{2}\right),$
where the number $\lambda$ satisfies
$\lambda^{2}=\frac{2}{\rho^{2}+1}.$ (21)
Because $Z$ is the dual of the self-dual form $\omega_{+}$, we have
${\langle}J_{x}v,w{\rangle}={\langle}{\langle}Z_{x},v\wedge w{\rangle}{\rangle}\;\;\text{for all}\;\,v,w\in T_{x}M.$
Using this relation together with (21), we obtain
$J_{x}e^{x}_{1}$ $=\frac{\lambda}{\sqrt{2}}\left(-\rho\,e^{x}_{2}+e^{x}_{3}\right),\;\;J_{x}e^{x}_{2}=\frac{\lambda}{\sqrt{2}}\left(\rho\,e^{x}_{1}-e^{x}_{4}\right),$
$J_{x}e^{x}_{3}$ $=-\frac{\lambda}{\sqrt{2}}\left(e^{x}_{1}+\rho\,e^{x}_{4}\right),\;\;J_{x}e^{x}_{4}=\frac{\lambda}{\sqrt{2}}\left(e^{x}_{2}+\rho\,e^{x}_{3}\right).$
From these expressions, and combining (3) with (ii2), we obtain (20). Hence (18) follows directly.
Therefore, the second fundamental form of the submanifold satisfies
@@
Immersions satisfying condition (22) have parallel second fundamental form (see Theorem 4 in [9]). Moreover, under the immersion $f$, each geodesic of $M^{4}$ is mapped into a plane circle. Such submanifolds were classified in [9] and [22]. From this classification, it follows that the submanifold is as described in part $(iii$-$b)$ of the theorem, thereby completing the proof for the case where the manifold is simply connected.